A student is skateboarding down a ramp that is 7.82 m long and inclined at 15.5° with respect to the horizontal. The initial speed of the skateboarder at the top of the ramp is 3.86 m/s. Neglect friction and find the speed at the bottom of the ramp.
Answer #1

here  

u = 3.86 m/s

s = 7.82 m

a = g*sin15.50 = 9.8*sin15.5 = 2.61 m/s2

Now applying the 3rd equation of NLM

v2 = u2 + 2*a*s

v2 = (3.86)2 + 2*2.61*7.82

v = 7.46 m/s